Two point charges are fixed on the y axis: a negative point charge q1 = -27 μC at y1 = +0.21 m and a positive point charge q2 at y2 = +0.37 m. A third point charge q = +8.8 μC is fixed at the origin. The net electrostatic force exerted on the charge q by the other two charges has a magnitude of 25 N and points in the +y direction. Determine the magnitude of q2.
1
Expert's answer
2020-09-21T08:33:25-0400

Explanations & Calculations


  • Force between q1 & q is attractive since those are opposite charges.
  • That force can be calculated to be,

"\qquad\qquad
\begin{aligned}
\small F_{q_1q} &= \small \frac{1}{4 \pi \epsilon_0} \times\frac{|q_1||q|}{(0.21m)^2}\\
&= \small 8.9875\times 10^{-3}\times \frac{27\times 8.8}{0.0441}\\
&= \small 48.4224 \,\,N
\end{aligned}"

  • Force between q2 & q is repulsive because of the like charges & that is,

"\qquad\qquad
\begin{aligned}
\small F_{q_2q} &= \small 8.9875\times 10^{-3}\times \frac{8.8\times q_2}{0.37^2}\\
&= \small 0.5777q_2 \,\,N
\end{aligned}"

  • Since the net force is directed along +y, "\small F_{q_1q}> F_{q_2q}" therefore,

"\qquad\qquad
\begin{aligned}
\small F_{q_1q}-F_{q_2q} &= \small 25 \,\,N\\
\small 48.4224-0.5777q_2 &= \small 25\\
\small q_2 &= \small \bold{40.5442}=\bold{40.5\mu C}
\end{aligned}"